# Evaluating Kubernetes Performance for GenAI Inference: From Automatic Speech Recognition to LLM Summarization

### v1

@@ Abstract
As Generative AI (GenAI), particularly inference, rapidly emerges as a dominant workload category, the Kubernetes ecosystem is proactively evolving to natively support its unique demands. This industry paper demonstrates how emerging Kubernetes-native projects can be combined to deliver the benefits of container orchestration, such as scalability and resource efficiency, to complex AI workflows. We implement and evaluate an illustrative, multi-stage use case consisting of automatic speech recognition and summarization. First, we address batch inference by using Kueue to manage jobs that transcribe audio files with Whisper models and Dynamic Accelerator Slicer (DAS) to increase parallel job execution. Second, we address a discrete online inference scenario by feeding the transcripts to a Large Language Model for summarization hosted using llm-d, a novel solution utilizing the recent developments around the Kubernetes Gateway API Inference Extension (GAIE) for optimized routing of inference requests. Our findings illustrate that these complementary components (Kueue, DAS, and GAIE) form a cohesive, high-performance platform, proving Kubernetes’ capability to serve as a unified foundation for demanding GenAI workloads: Kueue reduced total makespan by up to 15%; DAS shortened mean job completion time by 36%; and GAIE improved Time to First Token by 82%.
@@ 3.3.3 Distributed Inference
In the final stage of our workflow, we evaluate distributed inference performance for summarization tasks using the open-weight and publicly released Qwen3-8B model [23], deployed on Kubernetes through the llm-d stack. Specifically, we deploy eight replicas of vLLM, each serving an identical Qwen/Qwen3-8B model instance. The Qwen LLM family is developed by Alibaba, and we opted for this particular model due to its open-source nature and enterprise adoption. We first compare three configurations to quantify the overhead of GAIE: (i) a standard Kubernetes ClusterIP service for stateless load balancing, (ii) GAIE with only the InferenceGateway, and (iii) GAIE with both the InferenceGateway and InferencePool enabled, including its inference-scheduler. Additionally, we assess routing strategies within GAIE, contrasting a random scheduler with the precise-cache “well-lit path” [29] strategy that leverages KV-cache telemetry to route requests toward replicas with matching prefix tokens. An overview of the setup is shown in Figure 4.
@@ 4.2 Accelerator Slicing: DAS
At the time of this writing, the integration between the DAS operator and Kueue for GPU memory—aware scheduling of AI serving workloads was still in progress [37], so the following experiments were orchestrated directly through Kubernetes native jobs utilizing DAS. By exposing GPU memory as an extended resource, Kueue will be able to enforce quotas and admission control while DAS continues to manage MIG slice allocation. The integration will focus on long-running workloads (Pods, Deployments, StatefulSets) commonly used for inference and serving applications.
Before investigating efficiency gains, we wanted to gain a thorough understanding of the impact of running a Whisper transcription job on a MIG slice as opposed to a full GPU. To this end, we conducted an experiment with only eight total jobs, half of them running the medium model and half the large model. This setup ensured that waiting time was excluded from the job completion time measurement as both variants provided enough concurrency primitives: eight GPUs for no MIG setup and eight GPU slices for the MIG setup. Each configuration (no MIG vs. MIG) was run three times, the aggregated results are presented in Table 5.
@@
Increased execution parallelism also resulted in increased resource utilization. With DAS both the average tensor core utilization and the average DRAM activity were almost three times higher. Note that the moderate decrease of the average GPU utilization is due to the nature of MIG-slicing: with certain slice allocation patterns (e.g., two 3g.20gb GPU instance profiles scheduled on a single A100 GPU), the residual capacity may not be allocable, resulting in some streaming multiprocessors remaining idle and consequently lowering the overall utilization average. Despite this “slice padding”, the overall idle capacity had been significantly reduced. Also the average CPU utilization on the worker node was nearly four times higher with DAS (37%) compared to the not using DAS case (10%).
@@ 4.3 Distributed Inference: llm-d and Gateway API Inference Extension
To evaluate distributed inference performance using llm-d and the InferenceGateway (via the Gateway API Inference Extension, GAIE), we adopted the intelligent inference scheduling “well-lit path” in the final stage of our workflow. The setup deployed eight replicas of vLLM, each serving the Qwen/Qwen3-8B model, behind an OpenShift Service Mesh envoy proxy [40] acting as the GAIE gateway.
The first experiment investigates the latency overhead introduced by the GAIE components as explained in Section 3.3.3. To avoid potential latency introduced by the LLM runtime and accurately measure overhead, a single mocking replica of the vLLM Simulator [16] pod was used instead of a real vLLM instance. Figure 5 depicts the overhead where the x-axis is reflecting the concurrency level (the number of synchronous streams in parallel) and the y-axis the average Time to First Token (TTFT) in ms. Throughout the experiment, the ClusterIP and Inference Pool-disabled scenarios exhibit very similar TTFT. In contrast, in the Inference Pool-enabled scenario this latency is between 3 ms and 11 ms higher compared to the other two setups. TTFT increases with concurrency and the Inference Pool-enabled scenario exhibits the highest latency, especially at higher concurrency levels, introduced by additional routing overhead added by the gateway and endpoint picker components. This overhead is clearly offset by the latency improvements observed in the following results.
@@
To evaluate the routing strategies within GAIE, we compare two endpoint selection strategies: Random-Scheduling (Baseline) and Precise-Scheduling (“well-lit path”). The workload comprised the 32 previously transcribed earnings call audio clips, each averaging 8,500 tokens (totaling 272,000 tokens). Using GuideLLM, we generated 256 summarization requests (each of the 32 prompts repeated eight times) sent through the inference gateway, with a maximum output of 512 tokens per request for the summary. Each vLLM replica ran on an NVIDIA A100 40GB GPU, with a per-instance KV-cache capacity of 143,360 tokens, yielding an aggregate of 1,146,880 tokens across all GPUs. As the total prompt volume exceeds the per-GPU cache capacity, random routing inevitably incurs cache misses and LLM prefilling overhead. In contrast, Precise-Scheduling maximizes the reuse of hot prefixes across replicas, demonstrating measurable gains in cache hit rate, end-to-end latency, and overall throughput.
Figure 6 compares the TTFT between Random Scheduling and Precise-Scheduling. The x-axis displays the synchronous concurrent streams in parallel, while the y-axis, in logarithmic scale, represents the TTFT in milliseconds. On average, Random Scheduling ($\sim$500 ms) is approximately 6.25 times slower than Precise-Scheduling ($\sim$80 ms). Also, the variability is significantly higher for Random Scheduling, with a 95% confidence spread of 100 ms compared to about 6 ms for Precise-Scheduling. Additionally, the 99th percentile of TTFT for Random Scheduling is up to 20 times slower than Precise-Scheduling.
@@
Figure 7 illustrates the per-request end-to-end request latency difference between Random Scheduling and Precise-Scheduling routing strategies. The x-axis depicts the concurrency and the y-axis the latency. Random scheduling, on average, incurs a latency of approximately 1 second more compared to Precise-Scheduling. Moreover, for the 99th percentile of end-to-end latency, Random Scheduling can be up to 6 s slower. Additionally, it exhibits higher variability, as evident from the error bars and the significantly higher 99th percentile, suggesting that Precise-Scheduling provides better predictability.
@@ 4.4 Takeaways
Across all experiments, the results collectively demonstrate the complementary strengths of Kueue, DAS, and distributed inference scheduling (via GAIE) in improving the efficiency, predictability, and scalability of AI inference workloads on Kubernetes. Kueue provides deterministic, fair, and reproducible scheduling with negligible admission latency overhead, while priority and preemption mechanisms substantially reduce makespan (up to 15%) and improve throughput for large jobs. DAS enhances parallelism by dynamically partitioning GPUs into smaller slices, achieving up to a 36% reduction in mean job completion time despite modest per-job slowdowns from resource sharing. Finally, GAIE’s intelligent inference routing with Precise-Scheduling decreases latency, i.e., reducing TTFT by over 6 times and end-to-end latency by up to 6 s, while ensuring stable, predictable performance under high concurrency. Together, these findings highlight a clear design direction: combining queue-aware scheduling, dynamic accelerator partitioning, and intelligent inference routing yields an adaptive and efficient orchestration stack for large-scale, heterogeneous AI workloads, balancing throughput, fairness, and determinism across the stack.
@@ 6 Future Work
Future work will focus on extending the evaluation to encompass a broader set of AI workloads, including multimodal and streaming inference scenarios, to further validate the flexibility and scalability of the proposed workflow. As previously stated, native integration between Kueue and Dynamic Accelerator Slicing (DAS) is currently lacking, therefore we plan to have future work focusing on characterizing an end-to-end batch inference workflow that combines Kueue for workload queueing and DAS for accelerator slicing within Kubernetes. Additionally, we plan to integrate and assess emerging Kubernetes-native components for AI inference, such as gang scheduling and Dynamic Resource Allocation (DRA), to enhance the system’s efficiency and responsiveness under dynamic workloads. These efforts aim to generalize the presented framework and position it as a foundational blueprint for end-to-end AI workload orchestration on cloud-native infrastructure.

### v2

@@ Abstract
As Generative AI (GenAI), particularly inference, rapidly emerges as a dominant workload category, the Kubernetes ecosystem is proactively evolving to natively support its unique demands. This industry paper demonstrates how emerging Kubernetes-native projects can be combined to deliver the benefits of container orchestration, such as scalability and resource efficiency, to complex AI workflows. We implement and evaluate an illustrative, multi-stage use case consisting of automatic speech recognition and summarization. First, we address batch inference by using Kueue to manage jobs that transcribe audio files with Whisper models and Dynamic Accelerator Slicer (DAS) to increase parallel job execution. Second, we address a discrete online inference scenario by feeding the transcripts to a Large Language Model for summarization hosted using llm-d, a novel solution utilizing the recent developments around the Kubernetes Gateway API Inference Extension (GAIE) for optimized routing of inference requests. Our findings illustrate that these complementary components (Kueue, DAS, and GAIE) form a cohesive, high-performance platform, proving Kubernetes’ capability to serve as a unified foundation for demanding GenAI workloads: Kueue reduced total makespan by up to 15%; DAS shortened mean job completion time by 36%; and GAIE working in conjunction with llm-d improved tail Time to First Token latency by up to 90% even under high loads.
@@ 3.3.3 Distributed Inference
In the final stage of our workflow, we evaluate distributed inference performance for summarization tasks using the open-weight and publicly released Qwen3-8B model [23], deployed on Kubernetes through the llm-d stack. Specifically, we deploy eight replicas of vLLM, each serving an identical Qwen3-8B model instance. The Qwen LLM family is developed by Alibaba, and we opted for this particular model due to its open-source nature and enterprise adoption. We first compare three configurations to quantify the overhead of GAIE: (i) a standard Kubernetes ClusterIP service for stateless load balancing, (ii) GAIE with only the InferenceGateway, and (iii) GAIE with both the InferenceGateway and InferencePool enabled, including its inference-scheduler. Additionally, we assess routing strategies within GAIE, contrasting a random scheduler with the precise-cache “well-lit path” [29] strategy that leverages KV-cache telemetry to route requests toward replicas with matching prefix tokens. An overview of the setup is shown in Figure 4.
@@ 4.2 Accelerator Slicing: DAS
At the time of this writing, the integration between the DAS operator and Kueue for GPU memory aware scheduling of AI workloads was still in progress [37], so the following experiments were orchestrated directly through Kubernetes native jobs utilizing DAS. By exposing GPU memory as an extended resource, Kueue will be able to enforce quotas and admission control while DAS continues to manage MIG slice allocation. The integration will focus on long-running workloads (Pods, Deployments, StatefulSets) commonly used for inference and serving applications.
Before investigating efficiency gains, we wanted to gain a thorough understanding of the impact of running a Whisper transcription job on a MIG slice as opposed to a full GPU. To this end, we conducted an experiment with only eight total jobs, half running the medium model and half running the large model. This setup ensured that waiting time was excluded from the job completion time measurement as both variants provided enough concurrency primitives: eight GPUs for no MIG setup and eight GPU slices for the MIG setup. Each configuration (no MIG vs. MIG) was run three times, the aggregated results are presented in Table 5.
@@
Increased execution parallelism also resulted in increased resource utilization. With DAS, both the average tensor core utilization and the average DRAM activity were almost three times higher. Note that the moderate decrease of the average GPU utilization is due to the nature of MIG-slicing: with certain slice allocation patterns (e.g., two 3g.20gb GPU instance profiles scheduled on a single A100 GPU), the residual capacity may not be allocable, resulting in some streaming multiprocessors remaining idle and consequently lowering the overall utilization average. Despite this “slice padding”, the overall idle capacity had been significantly reduced. Also the average CPU utilization on the worker node was nearly four times higher with DAS (37%) compared to the not using DAS case (10%).
@@ 4.3 Distributed Inference: llm-d and Gateway API Inference Extension
To evaluate distributed inference performance using llm-d and the InferenceGateway (via the Gateway API Inference Extension, GAIE), we adopted intelligent inference scheduling through the Precise Prefix cache aware (also known as Precise Scheduling) “well-lit path"" in the final stage of our workflow. The setup deployed eight replicas of vLLM, each serving the Qwen/Qwen3-8B model, behind an OpenShift Service Mesh envoy proxy [40] acting as the GAIE gateway.
The first experiment investigates the latency overhead introduced by the GAIE components as explained in Section 3.3.3. To avoid potential latency introduced by the LLM runtime and accurately measure overhead, a single replica of the vLLM Simulator [16] pod was used instead of a real vLLM instance. Figure 5 depicts the overhead where the x-axis is reflecting the concurrency level (the number of synchronous streams in parallel) and the y-axis the average Time to First Token (TTFT) in ms. Throughout the experiment, the ClusterIP and Inference Pool-disabled scenarios exhibit very similar TTFT. In contrast, in the Inference Pool-enabled scenario this latency is between 3 ms and 11 ms higher compared to the other two setups. TTFT increases with concurrency and the Inference Pool-enabled scenario exhibits the highest latency, especially at higher concurrency levels, introduced by additional routing overhead added by the gateway and endpoint picker components. This overhead is clearly offset by the latency improvements observed in the results that follow with real-world LLM use cases.
@@
To evaluate the routing strategies within GAIE, we compare two endpoint selection strategies: Random Scheduling (Baseline) and Precise Scheduling (“well-lit path”). The workload comprised the 32 previously transcribed earnings call audio clips, each averaging 8,500 tokens (totaling 272,000 tokens). Using GuideLLM, we generated 256 summarization requests (each of the 32 prompts repeated eight times) sent through the inference gateway, with a maximum output of 512 tokens per request for the summary. Each vLLM replica ran on an NVIDIA A100 40GB GPU, with a per-instance KV-cache capacity of 143,360 tokens, yielding an aggregate of 1,146,880 tokens across all GPUs. As the total prompt volume exceeds the per-GPU cache capacity, random routing inevitably incurs cache misses and LLM prefilling overhead. In contrast, Precise Scheduling maximizes the reuse of hot prefixes across replicas, demonstrating measurable gains in cache hit rate, end-to-end latency, and overall throughput.
Figure 6 compares the TTFT between Random Scheduling and Precise Scheduling. The x-axis displays the synchronous concurrent streams in parallel, while the y-axis, in logarithmic scale, represents the TTFT in milliseconds. On average, Random Scheduling ($\sim$500 ms) is approximately 6.25 times slower than Precise Scheduling ($\sim$80 ms). Also, the variability is significantly higher for Random Scheduling, with a 95% confidence spread of 100 ms compared to about 6 ms for Precise Scheduling. Additionally, the 99th percentile of TTFT for Random Scheduling is up to 20 times slower than Precise Scheduling.
@@
Figure 7 illustrates the per-request end-to-end request latency difference between Random Scheduling and Precise Scheduling routing strategies. The x-axis depicts the concurrency and the y-axis the latency. Random scheduling, on average, incurs a latency of approximately 1 second more compared to Precise Scheduling. Moreover, for the 99th percentile of end-to-end latency, Random Scheduling can be up to 6 s slower. Additionally, it exhibits higher variability, as evident from the error bars and the significantly higher 99th percentile, suggesting that Precise Scheduling provides better predictability.
@@ 4.4 Takeaways
Across all experiments, the results collectively demonstrate the complementary strengths of Kueue, DAS, and distributed inference scheduling (via GAIE) in improving the efficiency, predictability, and scalability of AI inference workloads on Kubernetes. Kueue provides deterministic, fair, and reproducible scheduling with negligible admission latency overhead, while priority and preemption mechanisms substantially reduce makespan, by up to 15% and improve throughput for large jobs. DAS enhances parallelism by dynamically partitioning GPUs into smaller slices, achieving up to a 36% reduction in mean job completion time despite modest per-job slowdowns from resource sharing. Finally, GAIE’s intelligent inference routing with Precise Scheduling decreases latency, i.e., reducing 99th percentile TTFT by up to 90% and end-to-end latency by up to 25% , while ensuring stable, predictable performance even at the highest concurrency. Together, these findings highlight a clear design direction: combining queue-aware scheduling, dynamic accelerator partitioning, and intelligent inference routing yields an adaptive and efficient orchestration stack for large-scale, heterogeneous AI workloads, balancing throughput, fairness, and determinism across the stack.
@@ 6 Future Work
Future work will focus on extending the evaluation to encompass a broader set of AI workloads, including multimodal and streaming inference scenarios, to further validate the flexibility and scalability of the proposed workflow. As previously stated, native integration between Kueue and Dynamic Accelerator Slicing (DAS) is currently lacking; therefore we plan to have future work focusing on characterizing an end-to-end batch inference workflow that combines Kueue for workload queueing and DAS for accelerator slicing within Kubernetes. Additionally, we plan to integrate and assess emerging Kubernetes-native components for AI inference, such as gang scheduling and Dynamic Resource Allocation (DRA), to enhance the system’s efficiency and responsiveness under dynamic workloads. These efforts aim to generalize the presented framework and position it as a foundational blueprint for end-to-end AI workload orchestration on cloud-native infrastructure.